OpenStudy (anonymous):

can someone help me find the number of positive, negative and complex numbers of functions?

OpenStudy (anonymous):

the first is f(x) = x^6 +x^5 +x^4 + 4x^3 - 12x^2 +12

jimthompson5910 (jim_thompson5910):

in f(x), how many sign changes are there?

OpenStudy (anonymous):

hmmm. 2

jimthompson5910 (jim_thompson5910):

so there are at most 2 positive real roots you could have 1 positive real root or you could have no positive real roots

jimthompson5910 (jim_thompson5910):

now find f(-x) and tell me what you get

OpenStudy (anonymous):

f(-x)?

jimthompson5910 (jim_thompson5910):

replace EVERY x with '-x' to get f(x) = x^6 +x^5 +x^4 + 4x^3 - 12x^2 +12 f(-x) = (-x)^6 +(-x)^5 +(-x)^4 + 4(-x)^3 - 12(-x)^2 +12 I'll let you do the last step

OpenStudy (anonymous):

im still confused sorry

OpenStudy (anonymous):

-x is -1 ?

jimthompson5910 (jim_thompson5910):

simplify (-x)^6 to get ______ (fill in the blank)

OpenStudy (anonymous):

-6?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

remember that exponents mean that you multiply the number by itself that many number of times examples 7^3 = 7*7*7 8^5 = 8*8*8*8*8 9^14 = 9*9*9*9*9*9*9*9*9*9*9*9*9*9 ... note: there are 14 copies of 9 being multiplied I'm sure you get the idea

OpenStudy (anonymous):

yeah, but -x is not a number

OpenStudy (anonymous):

is it?

jimthompson5910 (jim_thompson5910):

yes it is, x stands in place of any number

jimthompson5910 (jim_thompson5910):

what is (-x)^2 equal to

OpenStudy (anonymous):

-2?

OpenStudy (anonymous):

no 2

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

how are you getting rid of x?

OpenStudy (anonymous):

x is 1 right

jimthompson5910 (jim_thompson5910):

how do you know that x = 1?

OpenStudy (anonymous):

i thought if there was no value in front it was 1

jimthompson5910 (jim_thompson5910):

you're thinking of x = 1*x

jimthompson5910 (jim_thompson5910):

but we don't know if x equals 1 or not

OpenStudy (anonymous):

so what number is x?

jimthompson5910 (jim_thompson5910):

we don't know

jimthompson5910 (jim_thompson5910):

which is why we just leave it as x

OpenStudy (anonymous):

so the function equals 12?

jimthompson5910 (jim_thompson5910):

how are you getting that?

OpenStudy (anonymous):

i don't have an idea

jimthompson5910 (jim_thompson5910):

(-x)^2 = (-x)*(-x) = ???

OpenStudy (anonymous):

-x^2

jimthompson5910 (jim_thompson5910):

(-x)^2 = (-x)*(-x) (-x)^2 = (-1*x)*(-1*x) (-x)^2 = ???

OpenStudy (anonymous):

i dont know

jimthompson5910 (jim_thompson5910):

-1 * -1 = ???

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

no, I'm multiplying

OpenStudy (anonymous):

1

jimthompson5910 (jim_thompson5910):

yes, and x*x = ???

OpenStudy (anonymous):

x

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

x^2

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

that means we have (-x)^2 = (-x)*(-x) (-x)^2 = (-1*x)*(-1*x) (-x)^2 = (-1*-1)*(x*x) (-x)^2 = (1)*(x^2) (-x)^2 = 1*x^2 (-x)^2 = x^2

jimthompson5910 (jim_thompson5910):

now I want you to tell me what (-x)^3 is equal to

OpenStudy (anonymous):

x^3

jimthompson5910 (jim_thompson5910):

you sure?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

do the steps I did for (-x)^2, but instead, you'll have (-x)^3 = (-x)*(-x)*(-x)

jimthompson5910 (jim_thompson5910):

(-x)^3 = (-x)*(-x)*(-x) = ???

OpenStudy (anonymous):

-x^3

jimthompson5910 (jim_thompson5910):

better

jimthompson5910 (jim_thompson5910):

(-x)^4 = ???

OpenStudy (anonymous):

x^4

jimthompson5910 (jim_thompson5910):

(-x)^5 = ???

OpenStudy (anonymous):

-x^5

jimthompson5910 (jim_thompson5910):

(-x)^6 = ???

OpenStudy (anonymous):

i get it if its odd its negative and even is positive

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

yeah that's a good way to remember it

jimthompson5910 (jim_thompson5910):

f(x) = x^6 +x^5 +x^4 + 4x^3 - 12x^2 +12 f(-x) = (-x)^6 +(-x)^5 +(-x)^4 + 4(-x)^3 - 12(-x)^2 +12 f(-x) = ???

OpenStudy (anonymous):

x^18+4?

jimthompson5910 (jim_thompson5910):

what have we been doing before I posted my last message?

OpenStudy (anonymous):

finding if they're negative or positive

jimthompson5910 (jim_thompson5910):

so what does f(-x) = (-x)^6 +(-x)^5 +(-x)^4 + 4(-x)^3 - 12(-x)^2 +12 simplify to?

OpenStudy (anonymous):

i don't know i get stuck after 4(-x)^3

jimthompson5910 (jim_thompson5910):

replace (-x)^3 with what it's equal to

OpenStudy (anonymous):

-x^3

jimthompson5910 (jim_thompson5910):

4(-x)^3 turns into 4(-x^3)

jimthompson5910 (jim_thompson5910):

so what does f(-x) = (-x)^6 +(-x)^5 +(-x)^4 + 4(-x)^3 - 12(-x)^2 +12 simplify to?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

i don't know man

jimthompson5910 (jim_thompson5910):

well we just went over what (-x)^6, (-x)^5, etc were equal to

jimthompson5910 (jim_thompson5910):

tell me what you think f(-x) = (-x)^6 +(-x)^5 +(-x)^4 + 4(-x)^3 - 12(-x)^2 +12 simplifies to?

OpenStudy (anonymous):

x^18+16

jimthompson5910 (jim_thompson5910):

how are you getting 18?

OpenStudy (anonymous):

12 + 4

jimthompson5910 (jim_thompson5910):

replace (-x)^6 with what it's equal to and you get what?

OpenStudy (anonymous):

x^6

jimthompson5910 (jim_thompson5910):

f(-x) = (-x)^6 +(-x)^5 +(-x)^4 + 4(-x)^3 - 12(-x)^2 f(-x) = x^6 +(-x)^5 +(-x)^4 + 4(-x)^3 - 12(-x)^2 Now do the same for (-x)^5

OpenStudy (anonymous):

-x^5

jimthompson5910 (jim_thompson5910):

f(-x) = (-x)^6 +(-x)^5 +(-x)^4 + 4(-x)^3 - 12(-x)^2 f(-x) = x^6 +(-x)^5 +(-x)^4 + 4(-x)^3 - 12(-x)^2 f(-x) = x^6 + (-x^5) +(-x)^4 + 4(-x)^3 - 12(-x)^2 f(-x) = x^6 - x^5 +(-x)^4 + 4(-x)^3 - 12(-x)^2 Now do the same for (-x)^4

OpenStudy (anonymous):

x^4

jimthompson5910 (jim_thompson5910):

f(-x) = (-x)^6 +(-x)^5 +(-x)^4 + 4(-x)^3 - 12(-x)^2 f(-x) = x^6 +(-x)^5 +(-x)^4 + 4(-x)^3 - 12(-x)^2 f(-x) = x^6 + (-x^5) +(-x)^4 + 4(-x)^3 - 12(-x)^2 f(-x) = x^6 - x^5 +(-x)^4 + 4(-x)^3 - 12(-x)^2 f(-x) = x^6 - x^5 + x^4 + 4(-x)^3 - 12(-x)^2 Now do the same for (-x)^3. See how this works?

OpenStudy (anonymous):

-x^3

jimthompson5910 (jim_thompson5910):

finish up and tell me what you get

OpenStudy (anonymous):

i don't know how to simplify this

jimthompson5910 (jim_thompson5910):

why do I get the feeling you just don't want to work

OpenStudy (anonymous):

? im trying

OpenStudy (anonymous):

x^5 + 8x?

jimthompson5910 (jim_thompson5910):

4(-x)^3 turns into ???

OpenStudy (anonymous):

-4x^3

jimthompson5910 (jim_thompson5910):

f(-x) = (-x)^6 +(-x)^5 +(-x)^4 + 4(-x)^3 - 12(-x)^2 f(-x) = x^6 +(-x)^5 +(-x)^4 + 4(-x)^3 - 12(-x)^2 f(-x) = x^6 + (-x^5) +(-x)^4 + 4(-x)^3 - 12(-x)^2 f(-x) = x^6 - x^5 +(-x)^4 + 4(-x)^3 - 12(-x)^2 f(-x) = x^6 - x^5 + x^4 + 4(-x)^3 - 12(-x)^2 f(-x) = x^6 - x^5 + x^4 + (-4x^3) - 12(-x)^2 f(-x) = x^6 - x^5 + x^4 - 4x^3 - 12(-x)^2 I'll let you do the last step

OpenStudy (anonymous):

12x^2

jimthompson5910 (jim_thompson5910):

f(-x) = (-x)^6 +(-x)^5 +(-x)^4 + 4(-x)^3 - 12(-x)^2 f(-x) = x^6 +(-x)^5 +(-x)^4 + 4(-x)^3 - 12(-x)^2 f(-x) = x^6 + (-x^5) +(-x)^4 + 4(-x)^3 - 12(-x)^2 f(-x) = x^6 - x^5 +(-x)^4 + 4(-x)^3 - 12(-x)^2 f(-x) = x^6 - x^5 + x^4 + 4(-x)^3 - 12(-x)^2 f(-x) = x^6 - x^5 + x^4 + (-4x^3) - 12(-x)^2 f(-x) = x^6 - x^5 + x^4 - 4x^3 - 12(-x)^2 f(-x) = x^6 - x^5 + x^4 - 4x^3 - 12x^2

OpenStudy (anonymous):

what do i do when i subtract x^6 and x^5. subtract exponents?